# Quantum Radiometric Calibration

### v1

@@ I.1 Photoelectric detection
where the voltages $U$ are transimpedance amplified photo currents measured on continuous-wave light. $U_{\mathrm{tot}}$ is the total measurement voltage and $U_{\mathrm{dark}}$ is the contribution without laser light. $U_{\mathrm{perf}}$ represents the perfect case, i.e. the unknown voltage that the measurement system would provide if every photon of the light beam (that is focused on the photodiode surface) would be converted into exactly one photo electron, combined with zero dark current. In the limit of negligible dark noise, $\eta_{\mathrm{QE}}$ approaches $\eta_{\mathrm{DE}}$. The challenge is to determine accurately $\eta_{\mathrm{DE}}$ and thus $\eta_{\mathrm{QE}}$, i.e. to calibrate a photodetector.
I.2 Concept
The basis of QRC is the measurement of the Heisenberg uncertainty product of the two orthogonal electric field quadratures $\hat{X}$ and $\hat{Y}$ of a (strongly) squeezed vacuum state using a BHD with a pair of photodiodes that are to be calibrated. The result is compared with the inferred uncertainty product of the measured state without any decoherence, which we call the ‘Heisenberg reference’ of our concept.
Squeezed states have an electric field uncertainty that is for a certain phase range below that of the corresponding ground state. They were first described in the early 1970s [15, 16, 17, 18, 19, 20], first experimentally realised in the mid 1980s [21, 22, 23, 24], and became a user application in 2011 for the first time when they improved the sensitivity of the GEO600 gravitational wave detector in a joint search for gravitational waves together with Virgo [25, 26].
@@ I.2 Concept
The lower bound is reached if the squeezed state is pure. The crucial fact is that the uncertainty product of a squeezed vacuum state increases successively with each lost photon! Using the setup’s total photon efficiency $\eta$, initial squeezed and anti-squeezed quadrature variances $\Delta\!^{2}\!\hat{X}^{\prime}$ and $\Delta\!^{2}\hat{Y}^{\prime}=1/\Delta\!^{2}\!\hat{X}^{\prime}$ change accordingly to
$\Delta\!^{2}\!\hat{X}$=\; $\eta\,\Delta\!^{2}\!\hat{X}^{\prime}+(1-\eta)\,$, (3)
$\Delta\!^{2}\hat{Y}$=\; $\eta\,\Delta\!^{2}\hat{Y}^{\prime}+(1-\eta)\,$,
where we use the fact that variances (or noise powers) of independent processes add up when superimposed [20]. The variances on the left sides in Eq. (3) can be precisely measured with the BHD allowing for the remaining two unknown quantities $\eta$ and $\Delta\!^{2}\!\hat{X}^{\prime}=1/\Delta\!^{2}\hat{Y}^{\prime}$ to be calculated. $\Delta\!^{2}\!\hat{X}^{\prime}$ represents the squeeze factor of the produced state inferred for zero decoherence. Removing this quantity from Eqs. (3) yields
@@
Fig. 1 illustrates the effect of an imperfect total quantum efficiency $\eta$ on the uncertainty product of a squeezed state according to Eqs. (3). The chosen squeeze factor and loss value match our experimental demonstration. The area in (a) represents the measured uncertainty product of a mixed, 10-dB squeezed state.
@@
According to Eq. 5, $\eta_{\mathrm{DE}}$ is found when $\eta$ and the other three efficiencies are determined. Here, $\eta_{\mathrm{esc}}$ is the photon escape efficiency from the squeezing resonator, $\eta_{\mathrm{prop}}$ the propagation efficiency from the same resonator to the measurement, and $\eta_{\mathrm{mm}}$ the modematching efficiency with respect to the local oscillator beam of the BHD. $\eta_{\mathrm{mm}}$ corresponds to the square of the interference visibility at the BHD beam splitter. It is optimized and measured while driving the phase $\theta$ between the LO and an auxiliary beam from the squeezing resonator. $\eta_{\mathrm{prop}}$ corresponds to the transmissivity of the optics along the path from the squeezing resonator to the photodiodes. It is measured with a single photo diode on the auxiliary beam at both ends of the optical path. The squeezing resonator’s photon escape efficiency $\eta_{\mathrm{esc}}$ could previously not be measured but just estimated and turned out to be the by far largest error contribution of the method [13]. The concept for measuring $\eta_{\mathrm{esc}}$ is a result of this work and will be discussed later. The basic principle for measuring $\eta$ in Eq. (5) will be further discussed here first.
I.3 Precision and accuracy of QCR
@@
The precision increases with an increasing absolute value of the derivative of the measured uncertainty product to $\eta$. The measured uncertainty product is according to Eq. (3)
$\!\!\!\!\Delta\!^{2}\!\hat{X}\cdot\Delta\!^{2}\hat{Y}$ $=\left[\eta\,\Delta\!^{2}\!\hat{X}^{\prime}+(1\!-\!\eta)\right]\!\cdot\!\left[\eta\,\Delta\!^{2}\hat{Y}^{\prime}+(1\!-\!\eta)\right]$
$=1+4(\eta-\eta^{2})\left[\frac{1}{4}(\Delta\!^{2}\!\hat{X}^{\prime}+\Delta\!^{2}\hat{Y}^{\prime})-\frac{1}{2}\right]$
$=1+4(\eta-\eta^{2})\langle\hat{n}\rangle\,$, (6)
@@
where $(\Delta\!^{2}\!\hat{X}^{\prime}+\Delta\!^{2}\hat{Y}^{\prime})/4-0.5=\langle\hat{n}\rangle$ is the expectation value for the number of quantum correlated photons in the ‘Heisenberg reference’ [20]. For the pure 13.2-dB squeezed state in Fig. 1(b) $\langle\hat{n}\rangle\approx 4.7$.
Here, we introduce the ‘precision scaling’ of QRC $|S_{P}|$, which includes the derivative of Eq. (6) to $\eta$ and the fact that the precision of the measurements of $\Delta\!^{2}\!\hat{X}$ and $\Delta\!^{2}\hat{Y}$ increases with detection efficiency.
$\!\!|S_{P}|$ $=\eta^{2}_{\mathrm{DE}}\!\left|\frac{d(1+4(\eta\!-\!\eta^{2})\langle\hat{n}\rangle)}{d\eta}\right|=\eta^{2}_{\mathrm{DE}}\,|4-8\eta|\langle\hat{n}\rangle\,$. (7)
@@
Thus, the QRC precision increases with $\langle\hat{n}\rangle$ (whereby a limit is set by the fact that the decoherence due to phase fluctuations must remain negligible, see method box) and with $\eta_{\mathrm{DE}}$, i.e. with the quality of the photodiode to be calibrated.
@@
The QRC approach offers a very high degree of accuracy, as potential systematic errors in independent measurements can be quantified and minimised to negligible values (see method box). All additional measurements are exclusively photoelectric in nature and take place in situ, i.e. in the same laser-optical QRC setup. Another advantage for accuracy is that QRC does not calibrate the direct current of the photodetectors, but rather the spectral power at the sideband frequency of the user application. QRC can provide a comprehensive calibration curve against sideband frequency.
——————————————————————— Method box – The most important test for ruling out systematic errors caused by the BHD is to verify its proportionality. The measured spectral noise power must be proportional to the LO power across the entire target measurement range when the signal input is blocked, i.e.  when the vacuum state is measured (see for instance Fig. 2 in Ref. [31]). If the noise power increases more rapidly, it is not the noise power of the ground state that is being measured, but that of other noise sources, e.g. scattered photons. If it increases more slowly, this is an indication of saturation of the amplifier electronics.
Regardless of the BHD, a systematic error can arise if the phase $\theta$ of the squeezed light fluctuates during the measurement [32]. This type of decoherence also leads to a reduction in the squeeze factor, even if all photons are detected. By operating the squeezing resonator close to its oscillation threshold, the phase noise becomes visible and can be quantified (see for instance Figs. 4 in Ref. [33]). Ideally, QRC uses a range of $\langle\hat{n}\rangle$ in which phase fluctuations are demonstrably irrelevant. ———————————————————————
I.4 QRC performed with 10-dB squeezed light
@@
We calibrated a pair of 1550-nm HQE photodiodes from Laser Components without protection window for s-polarized continuous-wave laser light at 1550 nm at a sideband frequency of 5 MHz under 20 angle of incidence. The semiconductor surface was free of dust particles. The experiment consisted of a resonator to generate squeezed vacuum states of light at 1550 nm and a BHD for their measurement using two of the photodiodes to be calibrated. The efficiencies $\eta_{\mathrm{esc}}$, $\eta_{\mathrm{prop}}$, and $\eta_{\mathrm{mm}}$ were maximized by selecting high-quality optics and optimizing the modematching aiming for a measured squeeze factor as high as possible. (The three efficiencies were quantified afterwards by separate measurements.) We used the most efficient approach to create squeezed light, which is resonator-enhanced degenerate parametric down-conversion (PDC) in a second-order non-linear crystal [22, 23, 24, 13, 34]. A schematic is shown in Fig. 2.
@@
The crystals in the SHG and squeezing resonators were periodically poled potassium titanyl phosphate (PPKTP). The crystal geometry was plane-convex and formed a half-monolithic resonator together with a coupling mirror. The crystal was coated to be highly reflective on the convex side and anti-reflective on the other for 1550 nm and 775 nm. The coupling mirror had a design reflectivity of $\approx\!83$ % for 1550 nm and $\approx\!97.5$ % for 775 nm light. The temperature of the crystal was stabilized to a setpoint that provided double-resonance as well as a good phase matching. The length of the squeezing resonator was stabilized using the Pound-Drever-Hall (PDH) technique [35]. The squeezed vacuum states at 1550 nm without any carrier light ‘escaped’ the squeezing resonator through the coupling mirror, were separated from the 775 nm light via a dichroic mirror, and superimposed with the local oscillator on the BHD’s beam splitter. We directly measured a 10-dB squeezed photon shot noise at a 775-nm pump power of slightly less than 11 mW. The anti-squeezing was close to 13 dB. Increasing the pump power increased the squeeze factor just marginally, while the anti-squeezing increased significantly. We used this measurement to estimate the root mean square (rms) phase noise [36] to $\Delta\theta\!<\!(2\!\pm\!1)$ mrad, which was negligible for the QRC result reported further down.
@@
To measure the propagation efficiency $\eta_{\mathrm{prop}}$, we once more blocked the second harmonic pump beam and coupled an auxiliary 1550-nm field through the PDC cavity from the under-coupled side. The transmitted light on cavity resonance was in the same mode as the squeezed vacuum and experienced the same optical losses along the down stream path. Measuring its power behind the coupling mirror $P_{\mathrm{cm}}$ and in front of each homodyne photodiode $P_{\mathrm{PD1,PD2}}$ allowed for determining the propagation efficiency $\eta_{\mathrm{prop}}=(P_{\mathrm{PD1}}+P_{\mathrm{PD2}})/P_{\mathrm{cm}}$. The transmission efficiency of the coupling mirror itself was dominated by its anti-reflective coating with AR = 500 ppm $\pm$ 500 ppm. We determined the propagation efficiency to be $\eta_{\mathrm{prop}}=(99.49\,\pm\,0.25)$%.
We measured the mode matching efficiency $\eta_{\mathrm{mm}}$ at the BHD beam splitter by using the auxiliary beam transmitted through the squeezing resonator, see Fig. 2 (modematching lenses and mirrors not shown). The differential phase $\theta$ was swept linearly by a piezo-actuated steering mirror and the resulting fringes were recorded using a single photodiode of the BHD. The fringe visibility was derived according to $v\!\!\;=\!\!\;(U_{\rm max}\!\!\;-\!\!\;U_{\rm min})/(U_{\rm max}\!\!\;+\!\!\;U_{\rm min})$, where $U_{\rm max,min}$ were the photoelectric voltages after subtraction of the voltage without any light. The mode matching efficiency $v^{2}$ was measured to $\eta_{\mathrm{mm}}=v^{2}=(99.11\,\pm\,0.17)\%$.
@@
The key measurement of QRC is that of the total quantum efficiency $\eta$. For this purpose, the variances of the squeezed and anti-squeezed quadratures in Eq. (6) were measured with the BHD. We repeatedly swept the BHD read-out phase $\theta$ to continuously rotate between measuring the maximally squeezed and the maximally anti-squeezed quadratures, see Fig. 4. This approach offers a few advantages over the alternative of locking the read-out phase to each individual quadrature. First of all, it reduces the complexity of the setup by omitting the stabilization of one degree of freedom. Additionally, slow drifts in the local oscillator and pump power can be neglected. Scanning the read-out phase also guarantees that there is no systematic error in the locking point and ensures that each quadrature is optimally measured at some point. The traces in Fig. 4 already include dark noise subtraction and normalisation to the ground state variance. The variance of the ground state was measured with the signal input of the BHD blocked. The variance of the dark noise was measured with all light fields blocked.
@@ II Result and Discussion
The two photodiodes that were calibrated in this work have an average detection efficiency for 1550 nm light of
$\eta_{\mathrm{DE}}=\,$ $\frac{(94.48\!\!\;\pm\!\!\;0.22)\%\,/(98.583\!\!\;\pm\!\!\;0.015)\%}{\!\!(99.49\!\!\;\pm\!\!\;0.25)\%\times(99.11\!\!\;\pm\!\!\;0.17)\%}$
$=\,$ $\,(97.20\!\!\;\pm\!\!\;0.37)\%.$ (12)
@@
Our error bar corresponds to the total combined standard deviation. It is dominated by statistical errors in the individual measurements and can be further reduced by increasing the measurement time, using lower-loss optics (in particular, better anti-reflective coatings on the optics) and also fewer optics for beam guidance and focusing on the photodiodes. This would then allow an increase in the squeeze value, which would further reduce the statistical error, see section I.3. Already with this experiment we almost achieve the combined standard deviation reported in Ref.[12]. Unlike [12], we exclusively use the photoelectric effect and not the pyroelectric effect, for which light must heat a material through absorption. The electrical noise powers produced are directly related to the lower bound of Heisenberg’s uncertainty relation, which is a unique selling point of QRS and has the potential to deliver significantly increased calibration accuracies in the future. Our value is smaller than that in [14]. The one-sigma standard deviations overlap if values without retro-reflection are used in both cases.
@@
We note that our previous work [13] did not distinguish between $\eta_{\mathrm{DE}}$ and $\eta_{\mathrm{QE}}$. The value of $(99.5\pm 0.5)$% given there corresponds to the detection efficiency and not the quantum efficiency defined according to Eq. (1). Generally, detection efficiency ($\eta_{\mathrm{DE}}$) is just an upper bound for quantum efficiency ($\eta_{\mathrm{QE}}$), see Eq. (1). In this work, we used an LO power of 10 mW per photodiode. Using the recorded dark noise power at 5 MHz and Eq. (1), we obtain the only slightly lower value of $\eta_{\mathrm{QE}}{\rm(10\,mW,5\,MHz)}={(96.9\!\!\;\pm\!\!\;0.4)}\%$.
III Summary and conclusions
The emerging field of quantum technologies requires unity-gain photodiodes with greater 99% quantum efficiency for continuous-wave light in the range of a few 10 mW. A prominent example is the planned European gravitational wave observatory, the Einstein Telescope (ET), which is intended to exploit 10-dB squeezed light at 1550 nm for its low-frequency detector. Another prominent example is the optical quantum computer based on the continuous variables $X$ and $Y$ [6]. . 1550 nm is the favoured wavelength [5, 7, 8]. Without high-efficiency photo diodes, the required fault tolerance cannot be achieved [8].
@@ III Summary and conclusions
We calibrated the presumably best photodiodes for 1550 nm on the market (Laser Components) to a detection efficiency of just $(97.20\!\!\;\pm\!\!\;0.37)\%$ (and a corresponding quantum efficiency of $(96.9\!\!\;\pm\!\!\;0.4)\%$). These values are insufficient for optical quantum computers and for the current design of the low-frequency detector of ET. The ball is now in the manufacturers’ court.
@@
The QRC method presented in detail and used here has several unique features compared to established methods. It uses only Heisenberg’s uncertainty principle and the photoelectric effect. No direct current (DC) signals are used for calibration, but rather noise power at relevant measurement frequencies. QRC is suitable for determining quantum efficiencies that depend on the measurement frequency. We have shown that absolute calibration with an uncertainty of only 0.37% is possible with squeezed vacuum states in the achievable 10 dB range. Squeeze lasers of this quality are commercially available [38]. The QRC method represents an important technology for the realisation of fault-tolerant optical CV quantum computers.
Acknowledgements.
@@
Author Contributions
L.A. performed the final experiments and the data analysis. J.-M.M. performed initial $\eta_{\rm esc}$-measurements. R.S. invented QRC and the $\eta_{\rm esc}$-measurement approach, developed the theoretical QRC description, and prepared the manuscript with the help of L.A.
Disclosures
The authors declare no conflicts of interest.
Data Availability Statement
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v2

@@ Abstract
where the voltages $U$ are transimpedance amplified photo currents measured on continuous-wave light. $U_{\mathrm{tot}}$ is the total measurement voltage and $U_{\mathrm{dark}}$ is the contribution without laser light. $U_{\mathrm{perf}}$ represents the perfect case, i.e., the unknown voltage that the measurement system would provide if every photon of the light beam (that is focused on the photodiode surface) would be converted into exactly one photo electron, combined with zero dark current. In the limit of negligible dark noise, $\eta_{\mathrm{QE}}$ approaches $\eta_{\mathrm{DE}}$. The challenge is to determine accurately $\eta_{\mathrm{DE}}$ and thus $\eta_{\mathrm{QE}}$, i.e., to calibrate a photodetector.
The basis of QRC is the measurement of the Heisenberg uncertainty product of the two orthogonal electric field quadratures $\hat{X}$ and $\hat{Y}$ of a (strongly) squeezed vacuum state using a BHD with a pair of photodiodes that are to be calibrated. The result is compared with the inferred uncertainty product of the squeezed state without any decoherence, which we call the ‘Heisenberg reference’. Squeezed states were first described in the early 1970s [15, 16, 17, 18, 19, 20], first experimentally realised in the mid 1980s [21, 22, 23, 24], and became a user application in 2011 for the first time when they improved the sensitivity of the GEO600 gravitational wave detector in a joint search for gravitational waves together with Virgo [25, 26].
@@
The lower bound is reached if the squeezed state is pure. The crucial fact is that the uncertainty product of a squeezed vacuum state increases successively with each lost photon, which is illustrated in Fig. 1. Using the setup’s total photon efficiency $\eta$, initial squeezed and anti-squeezed quadrature variances $\Delta\!^{2}\!\hat{X}^{\prime}$ and $\Delta\!^{2}\hat{Y}^{\prime}=1/\Delta\!^{2}\!\hat{X}^{\prime}$ change accordingly to [20] {align} Δ^2 ^X = η Δ^2 ^X’ + (1 - η) ,     Δ^2 ^Y = η Δ^2 ^Y’ + (1 - η)   . The variances on the left sides above can be precisely measured with the BHD allowing for the remaining two unknown quantities $\eta$ and $\Delta\!^{2}\!\hat{X}^{\prime}=1/\Delta\!^{2}\hat{Y}^{\prime}$ to be calculated. $\Delta\!^{2}\!\hat{X}^{\prime}$ represents the squeeze factor of the produced state inferred for zero decoherence. Removing this quantity from Eqs. (2) yields
@@
According to Eq. 4, $\eta_{\mathrm{DE}}$ is found when $\eta$ and the other three efficiencies are determined. Here, $\eta_{\mathrm{esc}}$ is the photon escape efficiency from the squeezing resonator, $\eta_{\mathrm{prop}}$ the propagation efficiency from the same resonator to the measurement, and $\eta_{\mathrm{mm}}$ the modematching efficiency with respect to the local oscillator beam of the BHD. $\eta_{\mathrm{mm}}$ corresponds to the square of the interference visibility at the BHD beam splitter. $\eta_{\mathrm{prop}}$ corresponds to the transmissivity of the optics along the path from the squeezing resonator to the photodiodes. It is measured with a single photo diode on the auxiliary beam at both ends of the optical path. The squeezing resonator’s photon escape efficiency $\eta_{\mathrm{esc}}$ could previously not be measured but just estimated and turned out to be the by far largest error contribution of the method [13]. The concept for measuring $\eta_{\mathrm{esc}}$ is a result of this work and will be discussed later. The basic principle for measuring $\eta$ in Eq. (4) will be further discussed here first.
@@
The precision increases with an increasing absolute value of the derivative of the measured uncertainty product to $\eta$. The measured uncertainty product is according to Eq. (2) {align} Δ^2 ^X ⋅Δ^2 ^Y = [ η Δ^2 ^X’ + (1-η) ] ⋅ [ η Δ^2 ^Y’ + (1-η) ]
= 1 + 4(η-η^2) [14 (Δ^2 ^X’ + Δ^2 ^Y’) - 12 ]
@@
= 1 + 4(η-η^2) ⟨^n ⟩  , where $(\Delta\!^{2}\!\hat{X}^{\prime}+\Delta\!^{2}\hat{Y}^{\prime})/4-0.5=\langle\hat{n}\rangle$ is the expectation value for the number of quantum correlated photons in the ‘Heisenberg reference’ [20]. For the pure 13.2-dB squeezed state in Fig. 1(b) $\langle\hat{n}\rangle\approx 4.7$.
@@
Here, we introduce the ‘precision scaling’ of QRC $|S_{P}|$, which includes the derivative of Eq. (1) to $\eta$ and the fact that the measurement precisions of $\Delta\!^{2}\!\hat{X}$ and $\Delta\!^{2}\hat{Y}$ are both proportional to the total efficiency $\eta$. {align} —S_P— = η^2  — d (1+ 4(η- η2) ⟨^n⟩)d η — = η^2  —4 - 8 η— ⟨^n ⟩  . Thus, the QRC precision increases with $\langle\hat{n}\rangle$ (whereby a limit is set by the fact that the decoherence due to phase fluctuations must remain negligible, see ‘end matter’) and with $\eta_{\mathrm{DE}}$, i.e., with the quality of the photodiode to be calibrated.
@@
The QRC approach offers a very high degree of accuracy, as potential systematic errors in independent measurements can be quantified and minimised to negligible values. All additional measurements are exclusively photoelectric in nature and take place in situ, i.e., in the same laser-optical QRC setup. Another advantage for accuracy is that QRC does not calibrate the direct current of the photodetectors, but rather the spectral power at the sideband frequency of the user application. QRC can provide a comprehensive calibration curve against sideband frequency.
@@
We calibrated at a sideband frequency of 5 MHz a pair of 1550-nm HQE photodiodes from Laser Components (without protection window) optimized for s-polarized continuous-wave laser light at 1550 nm under 20 angle of incidence. The semiconductor surface was free of dust particles. The experiment consisted of a resonator to generate squeezed vacuum states of light at 1550 nm and a BHD for their measurement using two of the photodiodes to be calibrated. The efficiencies $\eta_{\mathrm{esc}}$, $\eta_{\mathrm{prop}}$, and $\eta_{\mathrm{mm}}$ were maximized by selecting high-quality optics and optimizing the modematching aiming for a measured squeeze factor as high as possible. (The three efficiencies were quantified afterwards by separate measurements.) We used the most efficient approach to create squeezed light, which is resonator-enhanced degenerate parametric down-conversion (PDC) in a second-order non-linear crystal [22, 23, 24, 13, 31]. A schematic is shown in Fig. 2.
@@
The crystals in the SHG and squeezing resonators were periodically poled potassium titanyl phosphate (PPKTP). The crystal geometry was plane-convex and formed a half-monolithic resonator together with a coupling mirror. The crystal was coated to be highly reflective on the convex side and anti-reflective on the other for 1550 nm and 775 nm. The coupling mirror had a design reflectivity of $\approx\!83$ % for 1550 nm and $\approx\!97.5$ % for 775 nm light. The temperature profile of the crystal [32, 33, 34] was stabilized to a setpoint that provided double-resonance as well as a good phase matching. The length of the squeezing resonator was stabilized using the Pound-Drever-Hall (PDH) technique [35]. The squeezed vacuum states at 1550 nm without any carrier light ‘escaped’ the squeezing resonator through the coupling mirror, were separated from the 775 nm light via a dichroic mirror, and superimposed with the local oscillator on the BHD’s beam splitter. We directly measured a 10-dB squeezed photon shot noise at a 775-nm pump power of slightly less than 11 mW. The anti-squeezing was close to 13 dB. Increasing the pump power increased the squeeze factor just marginally, while the anti-squeezing increased significantly. We used this measurement to estimate the root mean square (rms) phase noise [36] to $\Delta\theta\!<\!(2\!\pm\!1)$ mrad, which was negligible for the QRC result reported further down.
@@
The propagation efficiency $\eta_{\mathrm{prop}}$ and the mode matching efficiency $\eta_{\mathrm{mm}}$ at the BHD beam splitter were measured in the established manner, which is described in the ‘end matter’ to this letter. The measured values yielded the $\eta_{\mathrm{prop}}=(99.49\,\pm\,0.25)$% and $\eta_{\mathrm{mm}}=v^{2}=(99.11\,\pm\,0.17)\%$.
@@
The key measurement of QRC is that of the total quantum efficiency $\eta$. For this purpose, the variances of the squeezed and anti-squeezed quadratures in Eq. (1) were measured with the BHD. We repeatedly swept the BHD read-out phase $\theta$ to continuously rotate between measuring the maximally squeezed and the maximally anti-squeezed quadratures, see Fig. 4. This approach offers a few advantages over the alternative of locking the read-out phase to each individual quadrature as used before [13, 14]. First of all, it reduces the complexity of the setup by omitting the stabilization of one degree of freedom. Additionally, slow drifts in the local oscillator and pump power can be neglected. Scanning the read-out phase also guarantees that there is no systematic error in the locking point and ensures that each quadrature is optimally measured at some point. The traces in Fig. 4 already include dark noise subtraction and normalisation to the ground state variance. The variance of the ground state was measured with the signal input of the BHD blocked. The variance of the dark noise was measured with all light fields blocked.
@@
The two photodiodes that were calibrated in this work have an average detection efficiency for 1550 nm light of {align} η_DE =  (94.48 ± 0.22)%  / (98.583 ± 0.015)%(99.49 ± 0.25)% ×(99.11 ± 0.17)%
@@
=  (97.20 ± 0.37)%. Our error bar corresponds to the total combined standard deviation. It is dominated by statistical errors in the individual measurements and can be further reduced by increasing the measurement time, using lower-loss optics and also fewer optics for beam guidance and focusing on the photodiodes. Already with this experiment we almost achieve the combined standard deviation reported in Ref.[12]. Unlike [12], we exclusively use the photoelectric effect and not the pyroelectric effect, for which light must heat a material through absorption.
@@
The quantum efficiency is derived from the detection efficiency, taking into account the measured dark noise at 5 MHz, see Eq. (1). For the local oscillator power of 10 mW used, we obtain the only slightly lower value of $\eta_{\mathrm{QE}}{\rm(10\,mW,5\,MHz)}={(96.9\!\!\;\pm\!\!\;0.4)}\%$. We note that in our previous work [13], we only specified the value for detection efficiency according to Eq. (1).
@@
In conclusion, we calibrated the presumably best photodiodes for 1550 nm on the market (Laser Components) to a detection efficiency of just $(97.20\!\!\;\pm\!\!\;0.37)\%$. This value is insufficient for optical quantum computers based on the continuous variables $X$ and $Y$ [5, 6, 7, 8] and for the current design of the low-frequency detector of the Einstein Telescope (ET), which is intended to exploit 10-dB squeezed light at 1550 nm. The ball is now in the manufacturers’ court.
@@
The QRC method presented in detail and used here has several unique features compared to established methods. It uses only Heisenberg’s uncertainty principle and the photoelectric effect. No direct current (DC) signals are used for calibration, but rather noise power at relevant measurement frequencies. QRC is suitable for determining quantum efficiencies that depend on the measurement frequency. We have shown that absolute calibration with an uncertainty of only 0.37% is possible with squeezed vacuum states in the achievable 10 dB range. Squeeze lasers of this quality are commercially available [38]. The QRC method represents an important tool for the realisation of fault-tolerant optical CV quantum computers.